# Magnetic Interaction in White Dwarf Binaries as Mechanism for Long-Period Radio Transients

### v1

@@ 3.2 Geometric configuration
Within this model, the particle flow that powers radio emission should be confined in the magnetic loop connecting the two stars. As a result, coherent radio emission can be only detected when the magnetic loop plane aligns with the line of sight, i.e. at conjunction (see Fig. 1 left panel for the top view of the orbit). In order to suppress emission from other phases, the radiating electrons are required to move relativistically. The fact that radio emission is observed when the MD is at the superior conjunction requires that the WD is an anti-parallel rotator, i.e. ${\bf\Omega\cdot\mu}<0$.
In the conjunction configuration alone cannot guarantee the detection of radio emission. Because of the relativistic motion of the particles, the emission beam is essentially along the tangential direction of the field lines. Radio emission can be observed only when the magnetic field line tangential direction directions are close to the observer. Figure 1 right shows the side view of the system, with a marginal case that the field tangential direction barely reaches the line of sight. When the orbital inclination angle $i$ is even smaller (e.g. close to the face on configuration), no coherent radio emission can reach the observer. Optical observations give the constraint that $i<74^{\circ}$ and $i>50^{\circ}$ from the observed radial velocity amplitude (de Ruiter et al., 2024). Our model provides another limit from the low end, with the minimum $i$ depending on the unknown angle between $\bf\Omega_{\star}$ and $\bf\mu$.
@@ 4 Radiation Mechanism
So far we have not specified the radiation mechanism. In this section, we survey various mechanisms and suggest that a relativistic version of electron cyclotron maser emission (ECME) is likely the emission mechanism.
@@ 4.1 Characteristic frequencies
$n_{\rm GJ}=$ $\frac{B_{\star}\Omega_{\star}}{2\pi qc}\left(\frac{r}{R_{\star}}\right)^{-3}\simeq(3.1\times 10^{-2}\ {\rm cm^{-3}})\ \xi\left(\frac{B_{\star}}{10^{7}\ \rm G}\right)$ (12)
×$\times\left(\frac{P}{125.5\ \rm min}\right)^{-1}\left(\frac{r}{10^{10}\ \rm cm}\right)^{-3}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3}.$
The plasma frequency corresponding to the GJ-density can be estimated as
$\omega_{p}=\sqrt{\frac{4\pi q^{2}n_{\rm GJ}}{\gamma m_{e}}}\simeq(9.9\times 10^{2}\ {\rm rad\ s^{-1}})\ \xi^{1/2}\left(\frac{B_{\star}}{10^{7}\ \rm G}\right)^{1/2}$ (13)
×$\times\left(\frac{P}{125.5\ \rm min}\right)^{-1/2}\left(\frac{r}{10^{10}\ \rm cm}\right)^{-3/2}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3/2}\left(\frac{\gamma}{10^{2}}\right)^{-1/2}.$
@@
$\omega_{B}=\frac{eB_{\star}}{m_{e}c}\left(\frac{R_{\star}}{r}\right)^{3}\simeq$ $(5.9\times 10^{10}\ {\rm rad\ s^{-1}})\ \left(\frac{B_{\star}}{10^{7}\ \rm G}\right)$ (14)
×$\times\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3}\left(\frac{r}{10^{10}\ \rm cm}\right)^{-3}.$
For typical parameters, one has $\omega_{B}\gg\omega\gg\omega_{p}$.
4.2 Disfavored mechanisms
We survey various emission mechanisms and can rule out the following mechanisms.
  • The required specific viewing geometry and the observed high linear polarization degrees suggest that the emitting electrons should be relativistic. This rules out the emission mechanisms invoking non-relativistic electrons, such as the non-relativistic version of the ECME.
  • Type III solar flares emit coherent radio emission via collective Langmuir wave plasma emission (Melrose, 2017). For a static plasma, the required plasma number density is $\sim 10^{7}\ \rm cm^{-3}$ in order to have $\omega_{p}$ reach the observed frequency at $\sim$100 MHz. This density is much greater than the GJ-density. So at least the non-relativistic plasma emission is not suitable to explain the observations. More complicated relativistic plasma emission may be possible, but it is unclear how an instability can be driven from such a system.
  • Relativistic electrons can radiate in the background magnetic field via synchrotron radiation or curvature radiation. The cooling timescale of synchrotron radiation can be estimated as
    $t_{\rm syn}$ $=\frac{6\pi m_{e}c}{\gamma\sigma_{\rm T}\beta^{2}B_{\rm bg}^{2}}\simeq(3.3\times 10^{5}\ {\rm s})\ \left(\frac{\gamma}{5}\right)^{-1}\left(\frac{B_{\star}}{10^{7}\ \rm G}\right)^{-2}$ (15)
    ×$\times\left(\frac{M}{0.8M_{\sun}}\right)^{2}\left(\frac{P}{125.5\ \rm min}\right)^{4}>P=7530\ {\rm s}.$
    One can see that because the $B$ field is not strong, the synchrotron cooling time is longer than the orbital period so that electrons cannot completely lose their perpendicular momenta. So curvature radiation is not relevant. For synchrotron radiation, the characteristic frequency $\omega_{\rm c,syn}\sim\gamma^{2}\omega_{B}$ is too high, because $\omega_{B}$ already exceeds the observed frequency for typical parameters (Eq.(14)). We therefore conclude that both synchrotron and curvature radiation are likely not the radiation mechanisms of these systems.
@@ 4.3 Relativistic electron cyclotron maser emission
We believe that a relativistic version of the ECME mechanism is likely operating in the system to power the observed coherent radio emission. Such a radiation emission mechanism (Twiss, 1958) has been widely accepted to interpret decametric emission from Jupiter and auroral kilometric radiation from Earth. The Jupiter emission is triggered by the similar unipolar induction mechanism due to Jupiter-Io interaction (Goldreich & Lynden-Bell, 1969). Three physical conditions are required to produce ECME: The first is population inversion of the energetic electrons in the direction perpendicular to the background magnetic field to achieve negative absorption, i.e. $\partial f/\partial p_{\perp}>0$, where $f$ is the particle distribution function and $p_{\perp}$ is the momentum perpendicular to the background magnetic field. The second condition requires $\omega_{B}\gg\omega_{p}$, which is satisfied here. Finally, one should satisfy the gyromagnetic resonance condition, which can be written as (Wu & Lee, 1979; Melrose, 2017)
$\omega-s\frac{\omega_{B}}{\gamma}-k_{\parallel}v_{\parallel}=0,\ s=0,\pm 1,\pm 2\dots,$ (16)
@@
where $s$ is the harmonic number, $k_{\parallel}$ and $v_{\parallel}$ denote the wave vector and velocity of the electron along the background magnetic field. We consider the case of $s=1$. The emission frequency can be estimated as
$\omega\simeq\frac{\omega_{B}}{\gamma}$ $\simeq(5.9\times 10^{8}\ {\rm MHz})\ \left(\frac{B_{\star}}{10^{7}\ \rm G}\right)\left(\frac{r}{10^{10}\ \rm cm}\right)^{-3}$ (17)
×$\times\left(\frac{\gamma}{10^{2}}\right)^{-1}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3},$
which is consistent with the typical frequency observed in ILT J1101 + 5521.
@@
The observed emission of ILT J1101+5521 is mostly linearly polarized but also carries circular polarization. The non-relativistic version of ECME gives dominant circularly polarized emission. Linear polarization can be observed only in the direction perpendicular to the magnetic field line. We invoke a relativistic version of the ECME mechanism with electrons moving with $\gamma\sim 100$. The direction of the dominant linear polarization is at $1/\gamma\sim 0.01$ with respect to the electron motion direction. The observation is consistent with the line of sight slightly deviates from the tangential direction of the magnetic field lines.
Since cyclotron emission has a characteristic frequency, one prediction of the ECME mechanism is that the spectrum is expected to be narrow if $\gamma$ is small (otherwise the spread of $\gamma$ would broaden the spectrum). A lower $\gamma$ also suggests a stronger circular polarization. We therefore predict an anti-correlation between the broadness of the spectrum and the degree of circular polarization $\Pi_{V}$. Future detection of LPRTs with narrow spectra and high $\Pi_{V}$ will bring direct support to our model.

### v2

@@ 3.2 Geometric configuration
Within this model, the particle flow that powers radio emission should be confined in the magnetic loop connecting the two stars. As a result, coherent radio emission can only be detected when the magnetic loop plane aligns with the line of sight, i.e. at conjunction (see Figure 1 left panel for the top view of the orbit). In order to suppress emission from other phases, the radiating electrons are required to move relativistically, consistent with the estimate in Eq. (10). The observation of radio emission when the MD is at superior conjunction suggests that electrons should move towards the WD direction with the parallel electric field direction pointing away from the WD. Such a configuration can be only achieved if the WD is an anti-parallel rotator, i.e. ${\bf\Omega\cdot\mu}<0$ (Ruderman & Sutherland, 1975).
The conjunction configuration alone cannot guarantee the detection of radio emission. Because of the relativistic motion of the particles, the emission beam is essentially along the tangential direction of the field lines. Radio emission can be observed only when the magnetic field line tangential directions are close to the observer. Figure 1 right shows the side view of the system, with a marginal case that the field tangential direction barely reaches the line of sight. When the orbital inclination angle $i$ is even smaller (e.g. close to the face on configuration), no coherent radio emission can reach the observer. For ITL J1101+5521, optical observations give the constraint of $i<74^{\circ}$ and $i>50^{\circ}$ from the observed radial velocity amplitude (de Ruiter et al., 2024). Our model provides another limit from the low end, with the minimum $i$ depending on the unknown angle between $\bf\Omega_{\star}$ and $\bf\mu$.
@@ 4 Radiation Mechanism
In this section, we investigate the relativistic ECME mechanism for coherent radio emission and the possible associated high energy counterparts including X-ray and $\gamma$-ray emission through relativistic boosting cyclotron radiation and inverse Compton scattering, respectively.
@@ 4.1 Characteristic frequencies
$n_{\rm GJ}=$ $\frac{B_{\star}\Omega_{\star}}{2\pi qc}\left(\frac{r}{R_{\star}}\right)^{-3}\simeq(1.6\times 10^{-4}\ {\rm cm^{-3}})\ \left(\frac{B_{\star}}{10^{6}\ \rm G}\right)$ (12)
×$\times\left(\frac{P}{125.5\ \rm min}\right)^{-1}\left(\frac{r}{2.67\times 10^{10}\ \rm cm}\right)^{-3}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3}.$
The total number density may be written as $n=\kappa n_{\rm GJ}$, where $\kappa$ is the multiplicity parameter. The plasma frequency corresponding to the GJ-density can be estimated as
$\omega_{p}=\sqrt{\frac{4\pi q^{2}n}{\gamma m_{e}}}\simeq(2.3\times 10^{2}\ {\rm rad\ s^{-1}})\ \kappa^{1/2}\left(\frac{B_{\star}}{10^{6}\ \rm G}\right)^{1/2}$ (13)
×$\times\left(\frac{P}{125.5\ \rm min}\right)^{-1/2}\left(\frac{r}{2.67\times 10^{10}\ \rm cm}\right)^{-3/2}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3/2}\left(\frac{\gamma}{10}\right)^{-1/2}.$
@@
$\omega_{B}=\frac{eB_{\star}}{m_{e}c}\left(\frac{R_{\star}}{r}\right)^{3}\simeq$ $(3.1\times 10^{8}\ {\rm rad\ s^{-1}})\ \left(\frac{B_{\star}}{10^{6}\ \rm G}\right)$ (14)
×$\times\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3}\left(\frac{r}{2.67\times 10^{10}\ \rm cm}\right)^{-3}.$
For typical parameters, one has $\omega>\omega_{B}\gg\omega_{p}$.
@@ 4.2 Relativistic electron cyclotron maser emission
We believe that a relativistic version of the ECME mechanism is likely operating in the system to power the observed coherent radio emission. An anisotropic momentum distribution of relativistic plasma might be generated due to synchrotron radiation cooling (Bilbao et al., 2024). The ECME mechanism (Twiss, 1958; Wu & Lee, 1979) has been widely accepted to interpret decametric emission from Jupiter, auroral kilometric radiation from Earth, and magnetic cataclysmic variables. The Jupiter emission is triggered by the similar unipolar induction mechanism due to Jupiter-Io interaction (Goldreich & Lynden-Bell, 1969). Three physical conditions are required to produce ECME: The first is population inversion of the energetic electrons in the direction perpendicular to the background magnetic field to achieve negative absorption, i.e. $\partial f/\partial p_{\perp}>0$, where $f$ is the particle distribution function and $p_{\perp}$ is the momentum perpendicular to the background magnetic field. The second condition requires $\omega_{B}\gg\omega_{p}$, which is satisfied here. Finally, one should satisfy the gyromagnetic resonance condition, which can be written as (Wu & Lee, 1979; Melrose, 2017)
$\omega$ $=\frac{s\omega_{B}}{\gamma\left(1-\frac{v_{\parallel}}{c}\cos\theta_{v}\right)}$ (15)
$=s\frac{\omega_{B}}{\gamma}+k_{\parallel}v_{\parallel},\ s=0,\pm 1,\pm 2\dots,$
@@
where $s$ is the harmonic number, $k_{\parallel}$ and $v_{\parallel}$ denote the wave vector and velocity of the electron along the background magnetic field. We consider the case of $s=1$. Thus the emission frequency of relativistic version of ECME can be estimated as
$\nu$ $=\delta\frac{\omega_{B}}{2\pi}\simeq\frac{\gamma\omega_{B}}{2\pi}\simeq(5.0\times 10^{8}\ {\rm Hz})$ (16)
×$\times\left(\frac{r}{2.67\times 10^{10}\ \rm cm}\right)^{-3}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3}\left(\frac{B_{\star}}{10^{6}\ \rm G}\right),$
where $\delta=1/[\gamma(1-\beta\cos\theta_{v})]$ is the Doppler factor, which $\simeq\gamma$ when $\theta_{v}\sim 1/\gamma$, $\beta=v_{\parallel}/c$ is the normalized velocity and its direction is along the background magnetic field. It should be pointed out that the magnetic field strength, Lorentz factor of the electron and viewing angle can significantly influence the emission frequency. One can see that for typical parameters, the emission frequency is consistent with the typical frequency of several hundred MHz as observed in LPRTs.
The observed emission of ILT J1101+5521 is mostly linearly polarized but also carries circular polarization. As shown in Figure 2, strong linear polarization with partial circular polarization is naturally expected for relativistic ECME. In Appendix A, we present a calculation of cyclotron radiation of one single particle in the comoving frame. We show that 100% linear polarization can be observed only in the direction perpendicular to the magnetic field line and circular polarization will show up in other viewing angles (see Eq. (A1)). For relativistic ECME, the brightest emission comes from the $1/\gamma$ cone, which corresponds to the $90^{\rm o}$ viewing angle in the comoving frame. After Lorentz transformation, 100% linear polarization can be observed when the line of sight is along the $1/\gamma$ viewing angle at the edge of the radiation cone in the lab frame (upper panel of Figure 2). More generally, when the line of sight is within the $1/\gamma$ cone (lower panel of Figure 2), mixed linear and circular polarization can be observed.
Figure 2: A cartoon picture for the relativistic ECME. The solid line denotes the line of sight. The main radiation cone is denoted as the orange cone with half opening angle $\sim 1/\gamma$.
@@
We invoke a relativistic version of the ECME mechanism with electrons moving with $\gamma\sim 10$ to account for the emission of ILT J1101+5521. This Lorentz factor is smaller than the maximum Lorentz factor achievable due to unipolar induction (Eq. (10)) so is reasonable. The direction of the dominant linear polarization is at $1/\gamma\sim 0.1$ with respect to the electron motion direction. The observation is consistent with the line of sight slightly deviates from the tangential direction of the magnetic field lines. The non-negligible circular polarization suggests that the line of sight is slightly inside the $1/\gamma$ cone.
The observed bandwidth of the brightest pulse in ILT J1101 + 5521 is $\sim 40\ \rm MHz$ which is extremely narrow but still not monochromatic. Since cyclotron emission has a characteristic frequency, one expects that the ECME mechanism carries a narrow spectrum. Relativistic ECME introduces a spectral broadening if the $\gamma$ distribution has a spread around a central value. The radiation flux drops significantly both at much higher or lower frequencies. This is consistent with observations.
4.3 High energy counterparts
Since relativistic particles are involved, this model also predicts high-energy emissions. In the following, we discuss possible high-energy radiation mechanisms and estimate the X-ray emission power. When electrons are accelerated along the background magnetic field and move towards the WD, both electron Lorentz factor and background magnetic field strength of the WD increase. However, since electrons are moving nearly along the field line with an extremely small pitch angle, synchrotron radiation cannot occur. In the rest frame without parallel motion, the electron is still non-relativistic and radiates cyclotron emission. The maximum observed radiation frequency is boosted by the maximum Lorentz factor (Eq. (10)), i.e.
$\omega_{\rm cyc}\simeq\gamma_{\rm max}\omega_{B}\simeq(3.2\times 10^{17}\ {\rm rad\ s^{-1}})\ \zeta\left(\frac{B_{\star}}{10^{6}\ \rm G}\right)^{2}\left(\frac{R_{c}}{0.2R_{\sun}}\right)$ (17)
×$\times\left(\frac{M}{0.8M_{\sun}}\right)^{-2/3}\left(\frac{P}{125.5\ \rm min}\right)^{-7/3}\left(\frac{r}{10^{9}\ \rm cm}\right)^{-3}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3},$
which is in the soft X-ray band. The X-ray emission is beamed, and usually deviate from the direction of radio emission (Figure 1).
The total number of observable emitting charged particles at any instant can be estimated as $N_{\rm tot}\sim\pi(l_{e}/\gamma)^{2}l_{e}n$, where $l_{e}$ is the typical length scale of the radiation region, noticing that the cross section of the emission beam is defined by $l_{e}/\gamma$ due to relativistic beaming. The received power is boosted by a factor of $\gamma^{2}$ since the observed time interval is shorter by a factor of $1-\beta\cos\theta_{v}\sim 1/\gamma^{2}$ with respect to the emission time interval. Consider that the electron moves non-relativistically in the perpendicular direction with normalized velocity $\beta\sim 0.5$, the maximum X-ray luminosity due to relativistically boosted cyclotron radiation can be calculated as
$L_{\rm X,cyc}^{\rm max}$ $\simeq\gamma_{\rm max}^{2}N_{\rm tot}r_{0}^{2}c\beta^{2}B_{\rm bg}^{2}$ (18)
$\simeq(1.3\times 10^{27}\ {\rm erg\ s^{-1}})\ \kappa\zeta\left(\frac{B_{\star}}{10^{6}\ \rm G}\right)^{5}\left(\frac{R_{c}}{0.2R_{\sun}}\right)$
×$\times\left(\frac{M}{0.8M_{\sun}}\right)^{-2/3}\left(\frac{P}{125.5\ \rm min}\right)^{-10/3}\left(\frac{l_{e}}{10^{10}\ \rm cm}\right)^{3}$
×$\times\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{9}\left(\frac{r}{10^{9}\ \rm cm}\right)^{-3}\left(\frac{\beta}{0.5}\right)^{2},$
where $r_{0}=e^{2}/m_{e}c^{2}$ is the classical electron radius.
Two LPRTs have X-rays detected. AR Scorpii with a 213.6-min period has an X-ray luminosity of $\sim 5\times 10^{30}\ {\rm erg\ s^{-1}}$. Such a luminosity is reachable if AR Scorpii WD has a stronger magnetic field (note the steep dependence $P_{\rm cyc}\propto B_{\star}^{5}$ in Eq. (18)). Recently, another LPRT DART/ASKAP J1832–0911 with a $44.2$-min period was detected in X-rays with a high luminosity of $\sim 10^{33}\ {\rm erg\ s^{-1}}$ (Wang et al., 2024). Such a high luminosity is difficult to reach within our model unless $B_{*}$ is extremely high. This source on the other hand has a supernova remnant association (Li et al., 2024), suggesting a different (e.g. young magnetar) origin.
X-rays and even $\gamma$-rays can also be produced via inverse Compton (IC) scattering of relativistic particles off the thermal photons from the stars. The surface temperature of a WD is $T=(10^{4}\ {\rm K})\ T_{4}$. At a distance $r$ from the WD, the photon energy density is $U_{\rm ph}=U_{\rm ph,\star}(r/R_{\star})^{-2}=a_{\rm rad}T^{4}(r/R_{\star})^{-2}$. The surface temperature of the WD with $T=10^{4}\ \rm K$ defines a peak photon energy $h\nu=2.82k_{B}T$ with corresponding peak frequency $\nu\simeq 5.9\times 10^{14}\ {\rm Hz}\ T_{4}$. The typical frequency of the IC photons can be estimated as
$\nu_{\rm ICS}\simeq\gamma^{2}\nu\simeq(5.9\times 10^{18}\ {\rm Hz})\ \left(\frac{\gamma}{10^{2}}\right)^{2}\left(\frac{T}{10^{4}\ \rm K}\right),$ (19)
which is in the X-ray band. Note that we have adopted a higher Lorentz factor $\gamma\sim 10^{2}$ than the one to account for radio emission ($\gamma\sim 10$). This is possible because electrons are continuously accelerated by the parallel electric field along the magnetic field line. Since magnetic field lines are curved, the X-ray emission, emitted at a different location, would be beamed in a somewhat different direction (see Figure 1) and can be detectable only under preferred viewing geometry.
The total observed X-ray IC luminosity can be estimated as
$L_{\rm X,IC}$ $\simeq\gamma^{2}N_{\rm tot}\gamma^{2}\sigma_{\rm T}cU_{\rm ph}$ (20)
$\simeq(7.1\times 10^{22}\ {\rm erg\ s^{-1}})\ \kappa\left(\frac{\gamma}{10^{2}}\right)^{2}\left(\frac{r}{10^{9}\ \rm cm}\right)^{-5}\left(\frac{T}{10^{4}\ \rm K}\right)^{4}$
×$\times\left(\frac{l_{e}}{10^{10}\ \rm cm}\right)^{3}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{5}\left(\frac{B_{\star}}{10^{6}\ \rm G}\right)\left(\frac{P}{125.5\ \rm min}\right)^{-1},$
where $\sigma_{\rm T}$ is the Thomson cross section and one can see that the observed luminosity is much smaller than that due to relativistically boosted cyclotron radiation (Eq. (18)).
As the electrons further accelerated toward the WD surface, Lorentz factor further increases. For the most extreme case, we assume that the electron Lorentz factor reaches the maximum ($\gamma\sim 10^{5}$) near the WD. The IC photon frequency becomes
$\nu_{\rm ICS}$ $\simeq\gamma^{2}\nu\simeq(5.9\times 10^{24}\ {\rm Hz})\ \left(\frac{\gamma}{10^{5}}\right)^{2}\left(\frac{T}{10^{4}\ \rm K}\right),$ (21)
which is in the $\gamma$-ray band. The corresponding observed $\gamma$-ray luminosity is
$L_{\rm\gamma,IC}$ $\simeq\gamma^{2}N_{\rm tot}\gamma^{2}\sigma_{\rm T}cU_{\rm ph}$ (22)
$\simeq(7.1\times 10^{28}\ {\rm erg\ s^{-1}})\ \kappa\left(\frac{\gamma}{10^{5}}\right)^{2}\left(\frac{r}{10^{9}\ \rm cm}\right)^{-5}\left(\frac{T}{10^{4}\ \rm K}\right)^{4}$
×$\times\left(\frac{l_{e}}{10^{10}\ \rm cm}\right)^{3}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{5}\left(\frac{B_{\star}}{10^{6}\ \rm G}\right)\left(\frac{P}{125.5\ \rm min}\right)^{-1}.$
The $\gamma$-ray flux is typically too low to be detected. Under special geometric configurations, $\gamma$-rays may be beamed to the WD and be reprocessed as thermal emission. Assuming that the thermal emission is re-emitted in a hot spot with radius $l\sim 0.01R_{*}$, the temperature is
$T_{\rm max,spot}$ $\simeq\left(\frac{L_{\rm\gamma,IC}}{\pi l^{2}\sigma}\right)^{1/4}\simeq(5.3\times 10^{4}\ {\rm K})$ (23)
×$\times\left(\frac{P_{\rm ICS}}{7\times 10^{28}\ {\rm erg\ s^{-1}}}\right)^{1/4}\left(\frac{l}{0.01R_{\star}}\right)^{-1/2},$
which is of the same order of the surface temperature of the WD. Here $\sigma$ is the Stefan–Boltzmann constant. If the heat is instead spread out to the full WD surface, the corresponding temperature is
$T_{\rm max,full}$ $\simeq\left(\frac{L_{\rm\gamma,IC}}{4\pi R_{\star}^{2}\sigma}\right)^{1/4}\simeq(3.8\times 10^{3}\ {\rm K})$ (24)
×$\times\left(\frac{P_{\rm ICS}}{7\times 10^{28}\ {\rm erg\ s^{-1}}}\right)^{1/4}\left(\frac{l}{0.01R_{\star}}\right)^{-1/2},$
which is below the WD surface temperature. In any case, the effect of this heating effect is likely undetectable.